The electric current I, is amperes, in a circuit varies directly as the voltage V. When 18 volts are applied, the current is 3 amperes. What is the current when 30 volts are applied?
Found 2 solutions by Boreal, Alan3354:
Answer by Boreal(15235)

V=kI
18=k*3
k=6 (which is the resistance, by the way)
30=6*I, so I=5 amps.

Answer by Alan3354(69443)

The electric current I, is amperes, in a circuit varies directly as the voltage V. When 18 volts are [sic] applied, the current is 3 amperes. What is the current when 30 volts are [sic] applied?
----------------
3*(30/18) = 5 amps
======================
volts is singular.
18 volts is applied ...